OpenStudy (anonymous):

Factor the trinomial: 2x2 + 7x + 3 A (2x - 1)(2x + 3) B (2x + 3)(x + 1) C (2x + 1)(x + 3) D (2x + 1)(2x + 3)

OpenStudy (pfenn1):

\[2x^2 +7x+3=(2x+a)(x+b)\]where \[ab=7\]\[a+b=3\]

OpenStudy (pfenn1):

Sorry, I did that backwards.\[a+b=7\]\[ab=3\]